Telescope Operator's Night Log for 24-25 Feb 2009 MST

Operator: J. Hill (in Tucson) assisted by J. Little (mtn)
Observers: V. Testa, F. Pedichini
Instrument Support: O. Kuhn (Tucson)

Summary

Worked about half the night through broken cirrus clouds. Lost the rest to weather. The winter chiller failed again, so we ran on the summer chiller, but it was also struggling. We were only able to run mirror ventilation about half the time.

Details (times in MST)

Open

17:20 Side vent doors open to cool dome facing east, roof in ventilation mode.

17:48 Open Shutter Doors facing east. Wind light from South, broken cirrus clouds, T=+5.5 degC, D=-15degC.

18:00 Enable DX Actuator #430

18:05 Polycom cuts off

18:22 Slew to 03:21 +20 for flats.

18:30 Taking sky flats

18:50 Winter Chiller CH0401 stops working. Stop mirror ventilation. Mirrors are +1.5 degC above ambient of 5 degC.

18:55 Switched to summer chiller on ECSGUI

18:59 Restart mirror ventilation with summer chiller.

19:08 +2000 Z7 to both mirrors

19:18 Stop mirror ventilation as summer chiller is losing ground. -3/-5 degC. But after turning off mirror ventilation, the temperature drop increased from -3 to -4 degc when the flow dropped from 515 to 350 L/min (the proportionality the is expected).

19:33 Sent Z4=+500 to Blue

19:39 Sent Z4=-1000 to Blue (clearly the wrong way)

19:40 Sent Z4=+1500 (in two steps) Seeing is ~1.5 arcsec.

19:43 start Left mirror ventilation as summer LWT has recovered to -8 degC.

19:49 start Right mirror ventilation also, still at LWT = -7.5 degC

20:00 Editted scale factor in play-ob.pl to get 32 sec fastextra exposures through clouds. (It is hardwired in the perl script for DOFPIA.)

Wayne called to say that he has food poisoning in Flagstaff, so healthy telescope operators are rare birds this week.

20:12 stop Right mirror ventilation to let chiller recover.

20:32 set encoder type to strip (the operator is rusty)

03:50:30 a clear jump of the blue guide star

20:56 start Right mirror ventilation (Left still running)

21:09 loosened astigmatism limit in dofpia to 400 nm rms.

Close

21:27 Close for clouds, and approaching radar echoes although they might be virga.

21:44 Switch back to winter chiller

21:50 Change directory of dofpia to the normal path of /home/lbcobs/LBCFPIA/lbcfpia/src/

21:55 Making some Play and Stop OB tests on LBC. Crashed lbc server while loading an OB after stopping a previous OB during telescope slew. Fernando says this is the first crash of the week, so I feel proud :).

22:03 Crashed again.

22:08 Fernando turns off LBC. lbcstar/lbckill

22:10 Winter chiller has failed again.

22:14 Crashed again loading the Lulin_V OB

22:17 Fernando loads a differnt OB successfully

22:21 Switch to the summer chiller with loop at -3.6 degC.

22:45 Made backup copies of LBCFPIA/lbcfpia and TCS/Configuration in /home/hill on abell.

Open

23:12 Heavy clouds have passed, so we reopen shutter doors and vent doors

23:14 Restart mirror ventilation on both sides with initial summer LWT at -8.4 degC.

23:20 sorting out permissions on the setup.csh file for IDL

23:28 Red filter error, had to restart dofpia.

23:28 Red pupils are remarkably elliptical in 062857! What causes this?

23:30 Summer chiller is only dropping 2 degC now, so it is losing ground. Flow is 574 L/min.

23:35 Reboot CMU since LBC-red is stuck in the flushing state.

23:36 Stop Right mirror ventilation to let chiller catch up. Now temperature drop is 3.25 for flow of 516. The relationship doesn't seem linear.

23:42 Set normal flow to 500 L/min rather than the normal 520 (reduced by 4%). This seems to make the temperature drop 3.30 degC (increased by 1.3%) but probably this is in the noise of the measurements.

23:50 Set normal flow to 450 L/min, Turn Right mirror ventilation on. But this forces bypass valves closed, so my experiment is foiled. Back to 520 L/min target.

23:54 Now getting 2.94 degC temperture drop with 555 L/min, so we are in the almost happy state again.

23:58 Enabled DX actuator #430. This is why we had a pile of astigmatism in the Red pupils. The actuator had tripped off when we were parked at horizon.

00:12 Blue guide star jumps

00:29 Having trouble with azimuth axis railing. Wind (LBTO) is 10 m/sec. Close Vent doors.

00:40 Stop both mirror ventilations to let the chiller recover.

01:00 Restart both mirror ventilations.

01:04 Open all vent doors halfway. Wind(LBTO) is 5-8 m/sec.

01:12 Humidity is rising - dewpoint up to -11 degC. Sky still OK, with clouds to South and East moving away.

01:15 Sent +1600 Z4 to Red/DX.

01:18 Chiller dropping 4.6 degC when flow was 400 L/min when MV had tripped off.

01:20 Sent +600 Z4 to Blue/SX.

01:25 Stopped both mirror ventilations as we seem to be heating the mirror rather than cooling when the chiller is above -6 degC LWT.

01:58 Blue guiding failed wierd CCD readout. 085818_1. It recovered after ~30 sec.

02:03 Sent +800 Z4 to Red.

02:07 Sent +800 Z4 to Red.

02:24 Blue guiding failed for the last minute of 092217. Clouds. Last frames were blank.

02:35 One test in clouds of the new dofpia copied from dofpia_jmh.pro.

Close

02:38 Close for clouds. This marks the end of this observing night.

-- JohnHill - 25 Feb 2009
Topic revision: r4 - 25 Feb 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback